OpenStudy (anonymous):

Write the product as a sum. 13 sin (x/2) cos (x/4) The answer I keep getting is 1/2 (sin(3x/4)+sin(x/4)) Is this wrong? or can it be simplified further as 1/2sinx??

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (inkyvoyd):

?

OpenStudy (inkyvoyd):

Sorry, I forget product-sum formulae as well @FoolForMath

OpenStudy (inkyvoyd):

@Callisto

OpenStudy (anonymous):

The product sum formula is sinUcosV= 1/2(sin(U+V)+(sin(U-V)))

OpenStudy (anonymous):

\[1/2[\sin(x/2+x/4)+\sin(x/2-x/4)]\]

OpenStudy (anonymous):

But the answer I get from that is incorrect??

OpenStudy (anonymous):

\[1/2[\sin(3x/4)+\sin(x/4)]\]

OpenStudy (anonymous):

You should multiply by 13

OpenStudy (anonymous):

what about 13?

OpenStudy (anonymous):

ohhhhhhh so 13/2 should be in the front? wow I can't believe I missed that

OpenStudy (anonymous):

The 13 that was in fron of the product in the question.

OpenStudy (anonymous):

yes thank you.

OpenStudy (anonymous):

With pleasure.

OpenStudy (anonymous):

FYI: There is cool name of this formula: http://mathworld.wolfram.com/ProsthaphaeresisFormulas.html